Your portfolio has $707 invested in Google, $255 invested in Apple, $414 invested in Facebook, and $640 invested in Cash.

What weight is Google in your portfolio? (answer to two decimal places)
Answer #1

Total value=(707+255+414+640)=$2016

Hence weight of Google=(707/2016)

=0.35(or 35.07% approx).